OpenStudy (anonymous):

Marie has 60 credit hours toward her graduation from college. Of the 60 hours, 1/3 are in her major (physics), 1/4 are in her minor (math), and the rest are electives. a. How many hours does Marie have in her major? b. How many hours does Marie have in her minor? c. How many hours are in core courses and electives?

OpenStudy (anonymous):

a. 1/3 * 60 = 60/3 = 20 hours b. 1/4 * 60 = 60/4 = 15 hours c. 60 - (20+15) = 25